OpenStudy (anonymous):

f(x)= 7/x^3-9/x^5. let F(x) be antiderivative of f(x) with F(1)=0 I got a c=1.25 but is wrong. need help, please

OpenStudy (anonymous):

I get F(x)= 9/4x^4-7/2x^2 +c

OpenStudy (anonymous):

For F(1) = 2.25-3.5+c=0. so C=1.25,right?

OpenStudy (anonymous):

i didnt reduce my stupid fraction

OpenStudy (anonymous):

of all the petty things i screw up! How do i give medals?

OpenStudy (anonymous):

where did ur reply go?

OpenStudy (anonymous):

\[F(x) = \frac{-7}{2x^2}+\frac{9}{4x^4}+c\] You were correct Then\[F(1) = c - \frac{5}{4}\] So c = 1.25 Nothing is wrong here

OpenStudy (anonymous):

So why did you think your answer is wrong?

OpenStudy (anonymous):

it doesnt check with webwork

OpenStudy (anonymous):

I believe you were correct, unless f(x) is not as same as you wrote :)

OpenStudy (anonymous):

its right. Thanks for the reassurance!

OpenStudy (anonymous):

How do i give medals to my tutors, like you?

OpenStudy (anonymous):

Press "Best Response" if you think my answer is deserved. Actually, I remove your doubts due to I love maths only, not because of loving medal